What does it mean to refer to a 20-in TV set or a 25-in TV set? Such units refer to the diagonal of the screen. A 15-in TV set also has a width of 12 inches. What is its height???

Answer by checkley77(12569)

12^2+H^2=15^2
144+H^2=225
H^2=225-144
H^2=81
H=SQRT81
H=9 INCHES.
PROOF:
12^2+9*2=15^2
144+81=225
225=225